# Tracking objects that change in appearance with phase synchrony

### v1

@@ 5 Discussion
Humans can track objects through the world even as they change in appearance, state, or visibility through occlusion. This ability underlies many everyday behaviors, from cooking to building, and as we have demonstrated, it remains an outstanding challenge for today’s machine vision systems. However, by taking inspiration from neuroscience, we have developed CV-RNN, a novel recurrent network architecture that implements attention through neural synchrony, which can do significantly better. Our CV-RNN performs significantly better than any other DNN tested on FeatureTracker, and rivaled or approached human accuracy and decision-making on all versions of the challenge.
Our findings are strongly related to Neuroscience work that suggests that phase synchronization is a key component of perceptual grouping. The binding-by-synchrony theory (Singer, 2007) suggests that when neurons synchronize their firing, the features they encode become bound to the same object. Thus, neural oscillations and the synchrony that they induce on neural populations may allow the brain to group features representing the same object in a visual scene (but see  (Roelfsema, 2023; Shadlen & Movshon, 1999) for alternative hypotheses). Our work does not necessarily favor binding-by-synchrony over other competing theories (Fries, 2015; Jensen et al., 2014) but is a proof-of-concept that neural synchrony helps object tracking.
By visualizing the strategies learned by the CV-RNN to solve FeatureTracker, we were able to generate a novel and testable hypothesis regarding how neural synchrony supports object tracking. If neural synchrony acts similarly in human brains as it does in the CV-RNN, we speculate that similar phase behavior could be found in LFP recordings. Neuronal populations encoding for the target should display a phase shift on incorrect trials as we see within the CV-RNN. We also expect that neurons encoding for a distractor will be inhibited during occlusions, which could lead to a decrease in power of recorded LFPs (see Figs. A.9 and A.15 to observe this in the CV-RNN).
@@ 6 Acknowledgments
Funding for this project was provided by the Office of Naval Research (N00014-19- 1-2029) and ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A0004). Additional support was provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
@@ A.1 Extended Discussion
Additionally, we observed that in several conditions, the models’ performance exceeded human accuracy. This behavior might be attributed to the simple statistics of the data, which the models can easily learn. While humans can effortlessly translate their strategies to naturalistic videos, the scalability of the CV-RNN remains untested. We leave this investigation for future work.
@@ A.2 Limitations
Another limitation of our work is that our CV-RNN is foundationally more computationally expensive than a standard and equally parametrized feedforward or recurrent neural network. This is because it uses complex-valued neurons in its attentional module. While we control the number of parameters between our CV-RNN and the RNN (InT) control (see Fig A.11.1), we currently lack an equivalent model in terms of the number of operations.
@@ A.11 Control experiments
Figure A.12: Extended benchmark including an RNN with the same number of parameters as CV-RNN (brown bar).
@@
We hypothesize that the CV-RNN’s failure to reach human performance in the OOD color conditions is due to a lack of prior knowledge of the entire colorspace. Because the models are trained on only half of the colorspace, the filters intended to represent the other half may not be initialized properly. Consequently, the model struggles to track the target under these conditions, not due to a failure of the circuit itself, but because the preprocessing layer (Conv3D in Table 4) does not provide accurate information to the circuit. To test this hypothesis, we train an RNN and a CV-RNN on a training distribution that includes objects exhibiting colors from the full colorspace and shapes evolving according to all the rules of the GoL combined. Once trained, we extract the weights of the preprocessing layer (Conv3D in Table 4), freeze them, and then train the RNN and CV-RNN circuits along with the classification layers. We then test the resulting models on all OOD conditions and compare them with other pre-trained models (pre-training on Kinetics400). We observe a significant improvement for both circuits (compared to the versions without pre-training), bringing the CV-RNN closer to human performance.

### v3

@@ 5 Discussion
Humans can track objects through the world even as they change in appearance, state, or visibility through occlusion. This ability underlies many everyday behaviors, from cooking to building, and as we have demonstrated, it remains an outstanding challenge for today’s machine vision systems. However, by taking inspiration from neuroscience, we have developed CV-RNN, a novel recurrent network architecture that implements attention through neural synchrony, which can do significantly better. Our CV-RNN performs significantly better than any other DNN tested on FeatureTracker, and rivaled or approached human accuracy and decision-making on all versions of the challenge. Our findings are strongly related to neuroscience work, which suggests that phase synchronization is a key component of perceptual grouping. The binding-by-synchrony theory (Singer, 2007) suggests that when neurons synchronize their firing, the features they encode become bound to the same object. Thus, neural oscillations and the synchrony that they induce on neural populations may allow the brain to group features representing the same object in a visual scene (but see  (Roelfsema, 2023; Shadlen & Movshon, 1999) for alternative hypotheses). Our work does not necessarily favor binding-by-synchrony over other competing theories (Fries, 2015; Jensen et al., 2014) but is a proof-of-concept that neural synchrony helps object tracking.
By visualizing the strategies learned by the CV-RNN to solve FeatureTracker, we were able to generate a novel and testable hypothesis regarding how neural synchrony supports object tracking. If neural synchrony acts similarly in human brains as it does in the CV-RNN, we speculate that similar phase behavior could be found in LFP recordings. Neuronal populations encoding for the target should display a phase shift on incorrect trials as we see within the CV-RNN. We also expect that neurons encoding for a distractor will be inhibited during occlusions, which could lead to a decrease in power of recorded LFPs (see Figs. A.9 and A.16 to observe this in the CV-RNN).
@@ 6 Acknowledgments
Our work is supported by ONR (N00014-24-1-2026), NSF (IIS-2402875) to T.S. and ERC (ERC Advanced GLOW No. 101096017) to R.V., as well as “OSCI-DEEP” [Joint Collaborative Research in Computational NeuroScience (CRCNS) Agence Nationale Recherche-National Science Fondation (ANR-NSF) Grant to R.V. (ANR-19-NEUC-0004) and T.S. (IIS-1912280)], and the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004) to R.V. and T.S. Additional support was provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
@@ A.1 Extended Discussion
Additionally, we observed that in several conditions, the models’ performance exceeded human accuracy. This behavior might be attributed to the simple statistics of the data, which the models can easily learn. While humans can effortlessly translate their strategies to naturalistic videos, the scalability of the CV-RNN remains untested. We leave this investigation for future work with for example benchmarks proposed by Cai et al. (2023).
@@ A.11 Control experiments
Figure A.12: Extended benchmark including an RNN (RNN-L) with the same number of parameters as CV-RNN (brown bar).
@@
We hypothesize that the CV-RNN’s failure to reach human performance in the OOD color conditions is due to a lack of prior knowledge of the entire colorspace. Because the models are trained on only half of the colorspace, the filters intended to represent the other half may not be initialized properly. Consequently, the model struggles to track the target under these conditions, not due to a failure of the circuit itself, but because the preprocessing layer (Conv3D in Table 4) does not provide accurate information to the circuit. To test this hypothesis, we train an RNN and a CV-RNN on a training distribution that includes objects exhibiting colors from the full colorspace and shapes evolving according to all the rules of the GoL combined. Once trained, we extract the weights of the preprocessing layer (Conv3D in Table 4), freeze them, and then train the RNN and CV-RNN circuits along with the classification layers. We then test the resulting models on all OOD conditions and compare them with other pre-trained models (pre-training on Kinetics400). We observe a significant improvement for both circuits (compared to the versions without pre-training), bringing the CV-RNN closer to human performance (see Fig. A.13).
@@
Including Self-Supervised models.
We include in Fig. A.14 a self-supervised model, VideoMae (ViT-S with patch size 16) to evaluate whether the visual representation of such model can help solve FeatureTracker. We use pre-trained weights on Kinetics400 and finetune the model on the training set of FeatureTracker. The model’s performance on the testing sets is very similar to the one of the supervised Transformers, suggesting a key role played by the architecture more than the training procedure to solve FeatureTracker.
We also trained DINO (Caron et al., 2021) on FeatureTracker to extract the features frame by frame (starting from pre-trained weights on ImageNet) and use an MLP or an RNN to perform the classification task. The model did not converge in training.
We further added results from the new model SAM2 (Ravi et al., 2024). Like DINO, this model is not designed for binary classification. Still, we can evaluate its ability to track the target across frames by comparing the position of the predicted mask with the actual position of the target. Specifically, we use the "tiny" pre-trained version of the model and initialize the first mask with the target’s position. We evaluate the ability of the model to keep track of the target as it changes in appearance by defining the overall accuracy as the number of videos where the predicted mask was close to the actual position of the target (IoU > 0.9). We perform this evaluation on 1,000 images taken from the in-distribution test set. We report an accuracy of $0.001875\%(\pm 7.65e-05)$.
We did not include these two models in Fig. A.14 as the bars would all be at $50\%$ in all the conditions.
Refer to captionRefer to caption(a)(b)
Figure A.14: Extended benchmark including a self-supervised model, Video-MAE (purple bar).
@@ A.12.1 Varying the trajectories of colors and positions
A.12.2 More naturalistic conditions.
We create two new versions of the datasets with non-static backgrounds and textured objects as a first step towards more naturalistic stimuli.
To generate a non-static background, we model the background as a 3D Perlin noise, starting from a random state. In practice, the background is now colored (each sample starts from a random color and does not overlap with the colors of the objects) and evolves over time with a temporal and spatial dynamic different from the one of the objects. Considering the texture condition, we generate 5 possible textures (checkerboard, stripes, dots, noise, none). Each object is assigned a texture that will remain constant over time (only the noise texture changes over time). We control spatial, shape, and color dynamics similarly to the original version and report the accuracy of the CV-RNN and the baselines in Figures  A.17 and  A.18.
Refer to captionRefer to caption(a)(b)
Figure A.17: Results of the CV-RNN (in orange) and the baselines on a version of FeatureTracker with non-static background.
Refer to captionRefer to caption(a)(b)
Figure A.18: Results of the CV-RNN (in orange) and the baselines on a version of FeatureTracker with textured objects.
A.12.3 Objects disappearing.
We finally evaluate whether the models can handle objects disappearing in the videos. To do that, we generate two new versions of the dataset, with objects moving in and out of the frame or containing objects that can vanish over time.
One version allows the objects to vanish by removing the last rule of the game of life (this rule being: “If no pixel is active, activate the pixel in the center”). The other version allows the objects to move along trajectories evolving in a window larger than the frame size. As a consequence, the objects can start their trajectories outside of the frame and appear in the frame at the middle of the video, or start inside the frame but move outside during the video and potentially reappear later. These new rules only apply to distractors and not to the target not to hamper the design of the task.
We first test the models on the respective test sets of these two versions using the models trained on the version where the objects never vanish or disappear. The results are shown in Fig. A.19, first row. While all the baselines show a clear drop in performance, the RNN and especially the CV-RNN’s behavior remain very consistent across conditions. As a control, we train the models on the version of the dataset with objects moving in and out and test them on the other conditions (see Fig. A.19, second row). This new training condition improves the performance on the test set with objects moving in and out but also improves the robustness on the other conditions. However, the CV-RNN remains overall more accurate on all versions of the task.
Figure A.19: The models learn the task where the videos contain a fixed number of objects always visible in the frames. We test their ability to keep track of the target when the distractors can move and out of the frame or vanish as their shape changes with time (first row). In the second row, we show the performance of models trained and tested with objects moving and out and tested on objects always in the frames or vanishing.